Minute of KAGRA commissioning meeting
- Date: 2018/3/6 11AM~12PM
- Attendance: Miyakawa, Uchiyama, Miyoki, Aso, Akutsu
- eZuce: Izumi, Enomoto, Kimura, Suzuki, Tomaru, Ushiba, Inoue, Hirose
1. Report of work of the last week and plan of work
- Alignment, ETMY: Enomoto, Izumi
- Yend baffle PD did not work, but power supply was off.
- Beam was found again at the ETMX and ETMY by alignment to IYA viewport using PR3. Beam will be returned to center this week.
- Beam at the REFL port will be checked this week, but beam will be limited by vacuum work, so time sharing is needed.
- We need absolutely beam monitor like camera, PD etc, even if all the chambers are closed.
- X end PD will be removed. Baffle PD will be located instead. Changing PD will be done around 3/26.
- Height of the PR2 and PR3 did not change much. Beam was misaligned mainly at the MC. QPD for MC transmitted monitor was recovered and the beam was aligned according to the MC monitor.
- LVDT had several jumps due to the power outage. The calibration also may change.
- Effect of Heat link to hardness of Yaw motion(Miyamoto, Ushiba)
- Current temperature is ~40K. Speed of lowering temperature increased in these 3 days.
2. Schedule of ETMX installtion
- See schedule
- Check list for ETMX, Center.
- Improved point for EMTX should be shared to Vacuum team. It will be listed and shared.
- Bug list is made by izumi-kun, it will be merged to above list.
3. Tasks or discussions
Feed back to BS for MICH
- Experimental side:
- Data analysis side:
- LSC model(Michimura, Izumi)
- Same as iKAGRA or full model?
- CAL model location? which PC?
Selection of storing DAQ channel -> DGS, before phase 1
- Check the total amount.
Screens at control room -> who, when?
- Needs laser safety information.
GW alert: online alert in the control room might be prepared. -> Togoshi
- Next meeting: 3/13